nightshade0773:

Every weekend, Trudy bikes to the park to play disc golf with her friends. Before leaving home, she fills her water bottle to the top. This weekend, Trudy drinks 3/8 of a liter on the way to the park. Now there is only 1/8 of a liter remaining in Trudy's water bottle. Use an equation to find the amount of water that Trudy's water bottle holds.

xxemilianaxx:

@smokeybrown @snowflake0531

nightshade0773:

great my brain will hurt

xxemilianaxx:

No lol smokeybrown and snow are good at explaining things and help you understand more

Answermyquestions:

there is 3/8 and there is 1/8 left so it hold 4/8 which = 1/2 of a liter

mhanifa:

You just need to add up the volumes

livigale27:

would it be 4/8 sorry i dont got the answer cuz im dumb but whatever (don't listen to me im guessing)

xxemilianaxx:

@livigale27 wrote:
would it be 4/8 sorry i dont got the answer cuz im dumb but whatever (don't listen to me im guessing)
Don't guess c:

snowflake0531:

x is the total amount of water in the water bottle \[x-\frac{3}{8} = \frac{1}{8}\] because the total amount of water subtracts by how much she drank, equals how much she has left \[x=\frac{4}{8}\] and then you can simplify that fraction to get the answer

livigale27:

umm okay

livigale27:

wait was i right then-

Answermyquestions:

so im right cool